# Shannon entropy of optimized proton-neutron pair condensates

### v1

@@ Abstract
Proton-neutron pairing and like-nucleon pairing are two different facets of atomic nuclear configurations. While like-nucleon pair condensates manifest their superfluidic nature in semi magic nuclei, it is not absolutely clear if there exists a T=0 proton-neutron pair condensate phase in $N=Z$ nuclei. With an explicit formalism of general pair condensates with good particle numbers, we optimize proton-neutron pair condensates for all $N=Z$ nuclei between 16O and 100Sn, given shell model effective interactions. As comparison, we also optimize like-nucleon pair condensates for their semi-magic isotones. Shannon entanglement entropy is a measurement of mixing among pair configurations, and can signal intrinsic phase transition. It turns out the like-nucleon pair condensates for semi-magic nuclei have large entropies signaling an entangled phase, but the proton-neutron pair condensates end up not far from a Hartree-Fock solution, with small entropy. With artificial pairing interaction strengths, we show that the general proton-neutron pair condensate can transit from an entangled T=1 phase to an entangled T=0 phase, i.e. pairing phase transition driven by external parameters. In the T=0 limit, the proton-neutron pair condensate turns out to be a purely P pair condensate with large entanglement entropy, although such cases may occur in cold atom systems, unlikely in atomic nuclei.
@@ I Introduction
Pairing is a fundamental feature of many-body systems ranging from superconducting electronic systems, ultracold atoms, to atomic nuclei. The atomic nuclei have an unique two-partite composition, as they are made of two types of fermions, i.e. protons and neutrons, and the interactions among them approximately respect isospin symmetry. T=1 pairing is symmetric for proton-proton, neutron-neutron and proton-neutron pairs, while T=0 pairing is only for proton-neutron pairs. The T=1 pairing of like-nucleons Bohr et al. (1958) has important influence in nuclear binding energies, isotopic abundances, moment of inertia, and semi-magic nuclear spectrum. In semi magic nuclei, the General Seniority scheme can successfully describe energetic spectrum and electromagnetic transition strengthsTalmi (1981), and it is believed that the ground state of a semi magic nucleus can be well approximated with a T=1 $S$-pair (angular momentum $J=0$) condensate, or the seniority-0 state. The breaking of $S$ pairs, or higher-seniority configurations, explain the low-lying excited states. The T=1 $S$-pair condensate has a superconducting nature similar to the BCS phase in condensed matter physics; if there exists a T=0 correlated proton-neutron pair condensate phase in the $N=Z$ nuclei, it would be an interesting feature of atomic nuclei the two-partite system.
@@
In Sec. II we introduce the formalism of projection after variation of a general proton-neutron pair condensate, as well as the definition of entanglement entropy we used in this work; in Sec. III we show the spectrum and B(E2) values from one proton-neutron pair condensate, and present systematic results of $N=Z$ nuclei from 18F to 98In, as well as pairing “phase” transition with artificial interactions; in Sec. IV we summarize this work and discuss the outlooks.
@@ II.1 Interactions and Configurations
where $|0\rangle$ is the bare fermion vacuum, and $\hat{c}^{\dagger}_{\alpha},\hat{c}^{\dagger}_{\beta}$ are fermion creation operators with $\alpha,\beta$ labelling single-particle states with quantum numbers $(nljm)$, which can be (though not restricted to) spherical harmonic bases of the shell model, and, finally, $A_{\alpha\beta}$ are the “pair structure coefficients”. Without losing generality, $A_{\alpha\beta}$ makes a skew complex matrix, though in practice $A_{\alpha\beta}$ is often restricted to be real. A Slater determinant of the Hartree-Fock method, a seniority-0 state, or a number projected component of the BCS/HFB vacuum are all special cases of the general pair condensate as in Eq. (4), therefore this trial wavefunction reserves more generalityRowe et al. (1991). In quantum chemical physics, a general pair condensate is also known as “geminal wavefunction” Moisset et al. (2022).
@@ II.2 Variation
Variational formulas of like-nucleon pair condensates has been proposed in different forms Rowe et al. (1991); Silvestre-Brac and Piepenbring (1982); Otsuka et al. (2001); Lei et al. (2020); Lu et al. (2022), and recently triaxially angular momentum projected states out of a like-nucleon pair condensate has been benchmarked in shell model spaces with effective interactionsLu et al. (2022). In this work we extend the formalism in Ref. Lu et al. (2022) from like-nucleon pair condensates to proton-neutron pair condensates. We assume a general proton-neutron pair condensate as the trial wavefunction ansatz,
$|\pi\nu;PC\rangle\equiv\frac{1}{n!}(\hat{A}^{\dagger}_{\pi\nu})^{n}|0\rangle,~{}~{}~{}~{}\hat{A}^{\dagger}_{\pi\nu}=\frac{1}{2}\sum_{\alpha\in\pi,\beta\in\nu}A_{\alpha\beta}\hat{c}^{\dagger}_{\alpha}\hat{c}^{\dagger}_{\beta},$ (5)
where $\alpha$ stands for proton orbits and $\beta$ stands for neutron orbits. The pair coefficients $A_{\alpha\beta}$ are parameters to be determined by variation, so to minimize the energy
One can restrict $A_{\alpha\beta}$ to good isospin a priori, by demanding
@@ II.3 Canonical basis and Shannon entropy
As the canonical basis in Eq.(14) is unique given the pair condensate (16), we use the Shannon entropy, though we believe using the invariant correlation entropy Volya and Zelevinsky (2003) leads to the same physical conclusions. We define the Shannon entropy as Guan et al. (2016)
@@
so to quantify how close a general pair condensate in Eq. (4) is to a Slater determinant, where $\{k_{1},k_{2},\cdots,k_{n}\}$ traverse all $C^{n}_{\Omega}$ different choices of $n$ different $(k,\bar{k})$ pairs out of all $\Omega$ ones, i.e. all the pairing configurations, and $p(k_{1}k_{2}\cdots k_{n})$ denotes the corresponding probability of the pair configurations. The logarithmic base $d$ equals $C_{\Omega}^{n}$, to scale $S$ such that $S\in[0,1]$ for any given particle number. In the limit of a Slater determinant, only one $p(k_{1}\cdots k_{n})$ value equals $1$ and others equal 00, and correspondingly the entropy $S=0$, signaling a mean field solution; in the limit of uniform pairing, $S$ reaches its maximum value $S=1$, signaling the “entangled phase”.
II.4 Triaxial angular momentum
The triaxial angular momentum projection via an linear algebraic approach Johnson and O’Mara (2017) is implemented, to restore good angular momentum of the proton-neutron pair condensate (5). When the maximum angular momentum is $10\sim 20\hbar$, the linear algebraic approach is about 10 times faster than the traditional quadrature method Johnson and Jiao (2019).
@@ III.1 Energetic spectrum and B(E2) values
Figure 5: Excitation energy for yrast states of 88Ru, 92Pd and 96Cd. “PN-PC” denotes yrast states projected from a proton-neutron pair condensate, “PP-NN-PC” denotes those projected from a proton-proton neutron-neutron pair condensate, “SM” stands for the full configuration-interaction shell model, and “EXP” stands for experimental data, “PHF” stands for Projected Hartree-Fock states. All the theoretical computations are in the $0p_{1/2}0p_{3/2}0f_{5/2}0g_{9/2}$ major shell with JUN45 effective interactions.
@@
While the vibrator limit may indicate that the $B(E2,I\rightarrow I-2)$ value increases linearly as the angular momentum $I$ of the yrast state increasesQi et al. (2011), shell model B(E2) values of these three nuclei are roughly constant. In Tab. 2 we present B(E2) values from the shell model, PN-PC and PP-NN-PC. The effective charges are taken as $e_{p}=1.5e,e_{n}=0.5e$, and the harmonic oscillator length is taken simply as $r_{0}=A^{1/6}fm$. Both PN-PC and PP-NN-PC generate B(E2) values qualitatively close to shell model results, albeit experimental B(E2) values are not available yet, as far as we know. Consistent with the excitation spectrum in Fig. 5, the PP-NN-PC results of B(E2) values are slightly better than the PN-PC.
@@ III.2 Shannon entropy for 18F $\sim$ 98In
In comparison, the entropy of semi-magic oxygen, calcium, nickel isotopes are almost all notably above 0.1, signaling a correlated neutron-neutron condensate for those nuclei. Therefore, the NN-PC results align with the conventional wisdom that there is a correlated, or superfluidic pair condensate in semi-magic nuclei. For the semi-magic nuclei, the error of the ground state energy of PVPC is about $0.5$ MeV, close to exact ground state of the shell model, in support for the T=1 correlated pair condensate “phase”. The PN-PC ground states are typically 1-2 MeV above the shell-model ground states, less accurate than the NN-PC. Considering such a deviation of PN-PC against shell-model, we note that these one-reference results may not be conclusive enough, albeit the absence of an entangled PN-PC hints negatively on the picture of an entangled proton-neutron pair condensate “phase”.
@@ III.3 Competition between T=1 and T=0 pair condensates
Both the entropy $S$ and the energetic descent $\Delta E$ are diagnostic signals of pair condensates based on comparison against Hartree-Fock solutions. It is shown that $S$ and $\Delta E$ are highly consistent, though not in the same units. When $x=0$ in the physical situation, the entropies and energies of PN-PC and PP-NN-PC are quite close, thus they construct dual descriptions. The PP-NN-PC energy descends a bit deeper, so the energy gain $\Delta E_{\rm PP-NN-PC}$ is larger than the PN-PC, though both of them have small entropies.
Figure 9: Fractions of different pairs with $J=0,\cdots,5$ and $T=0,1$ in the optimized pair of the proton-neutron pair condensate for 24Mg. When $x=0$ the physical solution has all kinds of pairs of T=0 and T=1 mixed up; when $x<0$ S and D pairs dominate; when $x>0$ the T=0 condensate dominates with PFH pairs mixed and then goes to a pure P pair condensate after $x>=0.8$. S, P, D, F, G, H stand for $J=0,1,2,3,4,5$ pairs respectively.
@@ III.4 Anatomy of the pairs
To look closer into the microscopic details of the proton-neutron pair condensates, we decompose them into components with good angular momentum and isospin,
$\hat{A}^{\dagger}_{\pi\nu}=\sum_{JMT}\hat{\mathbbm{A}}^{JMT\dagger}_{\pi\nu},$ (27)
where $\hat{\mathbbm{A}}^{JMT\dagger}_{\pi\nu}$ denotes a pair with good angular momentum quantum numbers $(J,M)$ and good isospin $T=0,1$. Norms of such pairs can be easily computed with generalized Wick’s theorem Jin-Quan et al. (1993), and the fraction of $(J,T)$ components in a given proton-neutron pair is defined as
In $1s0d$ shell the angular momentum of a pair can only be $J=(0,1,2,3,4,5)$, we label pairs with such angular momenta as (S,P,D,F,G,H) pairs. The fraction $f_{JT}$ is summed on $M=-J,\cdots,J$. In Fig. 9, we show the anatomy of the proton-neutron pairs studied in Fig. 8. In the realm of $x\in[-1,-0.2]$ the proton-neutron pair condensate consists of purely T=1 S, D pairs, as in conventional textbook knowledge Casten (1990). In the physical situation of $x=0$, T=0 and T=1 pairs are mixed up. We note that the stacked bars mean that both T=0 and T=1 pairs with a given $J$ can coexist, e.g. $J=2,3,4$ with $x=0$. More interesting is the two-stage evolution for $x>0$: when $x=0.2,0.4,0.6$ the $J$=odd T=0 pairs P, F, H dominate, and P pair weights increase as x increases; when $x=0.8,1.0$ (when the “phase” transition saturates) a purely P pair condensation appears. The onset of a purely P pair condensate under purely T=0 interactions confirm the situation explained by boson mapping Van Isacker et al. (2016). This also reveals that, the T=0 components under physical interactions are quite different from those under purely T=0 interactionsPoves and Martinez-Pinedo (1998). Only when the T=0 interactions dominates, does the PN-PC becomes a highly-correlated, P-pair-dominating condensate.
In conclusion, the proton-neutron pair condensate becomes purely T=0 or T=1 faster than the entropy implies in Fig. 8, $x=-0.2/0.2$ gives rise to purely T=1 or T=0 condensates. But the T=0 condensate may evolve further, from a PFH dominance to a purely P pair condensate, which requires a nearly pure T=0 Hamiltonian. In physical conditions, the T=0, J=3,5 pairs are in dominance, while the P pair dominance, onsets when the system approaches the pure T=0 force limit, which might occur in artificial cold atom systems, and seems unlikely in atomic nuclei.

### v2

@@ Abstract
Proton-neutron pairing and like-nucleon pairing are two different facets of atomic nuclear configurations. While like-nucleon pair condensates manifest their superfluidic nature in semi magic nuclei, it is not absolutely clear if there exists a T=0 proton-neutron pair condensate phase in $N=Z$ nuclei. With an explicit formalism of general pair condensates with good particle numbers, we optimize proton-neutron pair condensates for all $N=Z$ nuclei between 16O and 100Sn, given shell model effective interactions. As comparison, we also optimize like-nucleon pair condensates for their semi-magic isotones. Shannon entanglement entropy is a measurement of mixing among pair configurations, and can signal intrinsic phase transition. It turns out the like-nucleon pair condensates for semi-magic nuclei have large entropies signaling an entangled phase, but the proton-neutron pair condensates end up not far from a Hartree-Fock solution, with small entropy. With artificial pairing interaction strengths, we show that the general proton-neutron pair condensate can transit from an entangled T=1 phase to an entangled T=0 phase, i.e. pairing phase transition driven by external parameters. In the T=0 limit, the proton-neutron pair condensate optimized for 24Mg turns out to be a purely P pair condensate with large entanglement entropy, although such cases may occur in cold atom systems, unlikely in atomic nuclei.
@@ I Introduction
Pairing is a fundamental feature of many-body systems ranging from superconducting electronic systems, ultracold atoms, to atomic nuclei. The atomic nuclei have a unique two-partite composition, as they are made of two types of fermions, i.e. protons and neutrons, and the interactions among them approximately respect isospin symmetry. T=1 pairing is symmetric for proton-proton, neutron-neutron and proton-neutron pairs, while T=0 pairing is only for proton-neutron pairs. The T=1 pairing of like-nucleons Bohr et al. (1958) has important influence in nuclear binding energies, isotopic abundances, moment of inertia, and semi-magic nuclear spectrum. In semi magic nuclei, the general seniority scheme can successfully describe energetic spectrum and electromagnetic transition strengthsTalmi (1981), and it is believed that the ground state of a semi magic nucleus can be well approximated with a T=1 $S$-pair (angular momentum $J=0$) condensate, or the seniority-0 state. The breaking of $S$ pairs, or higher-seniority configurations, explain the low-lying excited states. The T=1 $S$-pair condensate has a superconducting nature similar to the BCS phase in condensed matter physics; if there exists a T=0 correlated proton-neutron pair condensate phase in the $N=Z$ nuclei, it would be an interesting feature of atomic nuclei the two-partite system.
@@
In Sec. II we introduce the formalism of projection after variation of a general proton-neutron pair condensate, as well as the definition of entanglement entropy we use in this work; in Sec. III we show the spectrum and B(E2) values from one proton-neutron pair condensate, and present systematic results of $N=Z$ nuclei from 18F to 98In, as well as pairing “phase” transition with artificial interactions; in Sec. IV we summarize this work and discuss the outlooks.
@@ II.1 Interactions and Configurations
where $|0\rangle$ is the bare fermion vacuum, and $\hat{c}^{\dagger}_{\alpha},\hat{c}^{\dagger}_{\beta}$ are fermion creation operators with $\alpha,\beta$ labelling single-particle states with quantum numbers $(nljm)$, which can be (though not restricted to) spherical harmonic bases of the shell model, and, finally, $A_{\alpha\beta}$ are the “pair structure coefficients”. Without losing generality, $A_{\alpha\beta}$ makes a skew complex matrix, though in practice $A_{\alpha\beta}$ is restricted to be real. A Slater determinant of the Hartree-Fock method, a seniority-0 state, or a number projected component of the BCS/HFB vacuum are all special cases of the general pair condensate as in Eq. (4), therefore this trial wavefunction reserves more generalityRowe et al. (1991). In quantum chemical physics, a general pair condensate is also known as “geminal wavefunction” Moisset et al. (2022).
@@ II.2 Variation
Variational formulas of like-nucleon pair condensates has been proposed in different forms Rowe et al. (1991); Silvestre-Brac and Piepenbring (1982); Otsuka et al. (2001); Lei et al. (2020); Lu et al. (2022), and recently triaxially angular momentum projected states out of a like-nucleon pair condensate has been benchmarked in shell model spaces with effective interactionsLu et al. (2022). In this work we extend the formalism in Ref. Lu et al. (2022) from like-nucleon pair condensates to proton-neutron pair condensates. We assume a general proton-neutron pair condensate as the trial wavefunction ansatz, $|\pi\nu;PC\rangle=(\frac{1}{2}\sum_{\alpha\in\pi,\beta\in\nu}A_{\alpha\beta}\hat{c}^{\dagger}_{\alpha}\hat{c}^{\dagger}_{\beta})^{n}|0\rangle$, as defined in Eq. (1). The pair coefficients $A_{\alpha\beta}$ are parameters to be determined by variation, so to minimize the energy $\min_{A_{\alpha\beta}}\langle\pi\nu;PC|\hat{H}|\pi\nu;PC\rangle/\langle\pi\nu;PC|\pi\nu;PC\rangle$. One can restrict $A_{\alpha\beta}$ to good isospin a priori, by demanding
@@ II.3 Canonical basis and Shannon entropy
As the canonical basis in (12) is unique given the pair condensate (14), we use the Shannon entropy, though we believe using the invariant correlation entropy Volya and Zelevinsky (2003) leads to the same physical conclusions. We define the Shannon entropy as Guan et al. (2016)
@@
where $p(k_{1}k_{2}\cdots k_{n})$ is the probability of finding the configuration $a^{\dagger}_{k_{1}}a^{\dagger}_{\bar{k}_{1}}a^{\dagger}_{k_{2}}a^{\dagger}_{\bar{k}_{2}}\cdots a^{\dagger}_{k_{n}}a^{\dagger}_{\bar{k}_{n}}$ in the pair condensate $|PC,n\rangle$,
$p(k_{1}k_{2}\cdots k_{n})=\frac{\nu^{2}_{k_{1}}\nu^{2}_{k_{2}}\cdots\nu^{2}_{k_{n}}}{\langle PC,n|PC,n\rangle},$ (17)
where $\nu_{k}$, as defined in (12), is the amplitude of non-collective pair $a^{\dagger}_{k}a^{\dagger}_{\bar{k}}$ in the collective pair $\hat{A}^{\dagger}$. The Shannon entropy $S$ quantifies how close a general pair condensate in Eq. (4) is to a Slater determinant, where $\{k_{1},k_{2},\cdots,k_{n}\}$ traverse all $C^{n}_{\Omega}$ different choices of $n$ different $(k,\bar{k})$ pairs out of all $\Omega$ ones, i.e. all the pairing configurations. The logarithmic base $d$ equals $C_{\Omega}^{n}$, to scale $S$ such that $S\in[0,1]$ for any given particle number. In the limit of a Slater determinant, only one $p(k_{1}\cdots k_{n})$ value equals $1$ and others equal 00, and correspondingly the entropy $S=0$, signaling a mean field solution; in the limit of uniform pairing, $S$ reaches its maximum value $S=1$, signaling the “entangled phase”.
II.4 Triaxial angular momentum projection
The triaxial angular momentum projection via an linear algebraic approach Johnson and O’Mara (2017) is implemented, to restore good angular momentum of the proton-neutron pair condensate (LABEL:pn-pc). When the maximum angular momentum is $10\sim 20\hbar$, the linear algebraic approach is about 10 times faster than the traditional quadrature method Johnson and Jiao (2019).
@@ III.1 Energetic spectrum and B(E2) values
Figure 5: Excitation energy for yrast states of 88Ru, 92Pd and 96Cd. “PN-PC” denotes yrast states projected from a proton-neutron pair condensate, “PP-NN-PC” denotes those projected from a proton-proton neutron-neutron pair condensate, “SM” stands for the full configuration-interaction shell model, and “EXP” stands for experimental data, “PHF” stands for Projected Hartree-Fock states. All the theoretical computations are in the $1p_{1/2}1p_{3/2}0f_{5/2}0g_{9/2}$ major shell with JUN45 effective interactions.
@@
While the vibrator limit may indicate that the $B(E2,I\rightarrow I-2)$ value increases linearly as the angular momentum $I$ of the yrast state increasesQi et al. (2011), shell model B(E2) values of these three nuclei are roughly constant. In Tab. 2 we present B(E2) values from the shell model, PN-PC and PP-NN-PC. The effective charges are taken as $e_{p}=1.5e,e_{n}=0.5e$, and the harmonic oscillator length is taken simply as $r_{0}=A^{1/6}fm$. Both PN-PC and PP-NN-PC generate B(E2) values qualitatively close to shell model results, except for B(E2, $8^{+}_{1}\rightarrow 6^{+}_{1}$) and B(E2, $10^{+}_{1}\rightarrow 8^{+}_{1}$) in 96Cd, albeit experimental B(E2) values are not available yet, as far as we know. Consistent with the excitation spectrum in Fig. 5, the PP-NN-PC results of B(E2) values are slightly better than the PN-PC.
@@ III.2 Shannon entropy for 18F $\sim$ 98In
In comparison, the entropy of semi-magic oxygen, calcium, nickel isotopes are almost all notably above 0.1, signaling a correlated neutron-neutron condensate for those nuclei. Therefore, the NN-PC results align with the conventional wisdom that there is a correlated, or superfluidic pair condensate in semi-magic nuclei. For the semi-magic nuclei, the error of the ground state energy of PVPC is about $0.5$ MeV, close to exact ground state of the shell model, in support for the T=1 correlated pair condensate “phase”. The PN-PC ground states are typically 1-2 MeV above the shell-model ground states, less accurate than the NN-PC. Considering such a deviation of PN-PC against shell-model, we note that these one-reference results may not be conclusive enough, albeit the absence of an entangled PN-PC hints negatively against the picture of an entangled proton-neutron pair condensate “phase”.
@@ III.3 Competition between T=1 and T=0 pair condensates
Both the entropy $S$ and the energetic descent $\Delta E$ are diagnostic signals of pair condensates based on comparison against Hartree-Fock solutions. It is shown that $S$ and $\Delta E$ are highly consistent. When $x=0$ in the physical situation, the entropies and energies of PN-PC and PP-NN-PC are quite close, thus they construct dual descriptions. The PP-NN-PC energy descends a bit deeper, so the energy gain $\Delta E_{\rm PP-NN-PC}$ is larger than the PN-PC, though both of them have small entropies.
Figure 9: Fractions of different pairs with $J=0,\cdots,5$ and $T=0,1$ in the optimized pair of the proton-neutron pair condensate for 24Mg, given the artificial Hamiltonian in (23). When $x=0$ the physical solution has all kinds of pairs of T=0 and T=1 mixed up; when $x=-1$, T=1 S and D pairs dominate; when $x=1$ the T=0 P pairs dominate. S, P, D, F, G, H stand for $J=0,1,2,3,4,5$ pairs respectively.
@@ III.4 Anatomy of the pairs
To look closer into the microscopic details of the proton-neutron pair condensates, we decompose them into components with good angular momentum and isospin, $\hat{A}^{\dagger}_{\pi\nu}=\sum_{JMT}\hat{\mathbbm{A}}^{JMT\dagger}_{\pi\nu}$, as mentioned in Sec. II.2. Norms of such pairs can be easily computed with generalized Wick’s theorem Jin-Quan et al. (1993), and the fraction of $(J,T)$ components in a given proton-neutron pair, $f_{JT}$, is defined in (10).
In $1s0d$ shell the angular momentum of a pair can only be $J=(0,1,2,3,4,5)$, we label pairs with such angular momenta as (S,P,D,F,G,H) pairs. In Fig. 9, we show the anatomy of the proton-neutron pairs studied in Fig. 8, with $x=0,\pm 1$, corresponding to physical interactions and artificial interactions in the T=1, T=0 limits. When the external parameter in (23), $x=-1$, i.e. in the T=1 limit, the optimized proton-neutron pair condensate is made of T=1 S, D pairs. This is in alignment with the conventional textbook knowledge that T=1 low-angular-momentum pairs are important Casten (1990). In the physical situation of $x=0$, the optimized proton-neutron pair condensate is a mixture of both T=1 and T=0 pairs. The stacked bars in Fig. 9 mean that both T=0 and T=1 pairs with a given $J$ can coexist, e.g. there are both T=0 and T=1 D, F, G pairs when $x=0$. In the T=0 limit, the fraction of T=0 P pair increases, and there onsets a pure P-pair condensate Van Isacker et al. (2016). This also reveals that, the T=0 components under physical interactions are quite different from those under purely T=0 interactionsPoves and Martinez-Pinedo (1998). Only when the T=0 interactions dominates, does the PN-PC becomes a highly-entangled (with large entropy), P-pair-dominating condensate.
With angular momentum projection techniques, we find that four proton-neutron P-pairs in 24Mg are in alignment, coupling into total angular momentum $J=4$. Similar numerical experiments show that, in the T=0 limit of the modified USDB and GX1A interactions respectively, 32S, 48Cr, 72Kr all end up with a purely P-pair condensate. However, curiously, in the $1p0f_{5/2}0g_{9/2}$ with the JUN45 interactions, in the T=0 limit, 64Ge ends up a P-pair condensate, but 92Pd ends up a $J=9$ pair condensate.